OpenStudy (anonymous):

integrate tan^2xdx=?????

OpenStudy (jamesj):

\[ 1 + \tan^2 x = \sec^2 x \] Hence \[ \int \tan^2 x dx = \int (\sec^2 x - 1) dx \] = ...

OpenStudy (anonymous):

tan^2(x) = sec^2(x) - 1 Sub u = tan(x) and du = sec^2(x)dx

OpenStudy (anonymous):

tnx :)

OpenStudy (anonymous):

so on?